# Near-Optimal Approximation of Matrix Functions by Lanczos-FAN. Amsel, T. Chen, A. Greenbaum, C. Musco, and C. Musco

### v2

@@ 2.2 Notation
Note that $m_{1,q}=m$, so $r_{q}=r(x)$. Recall that for any function $f$, the Lanczos-FA approximation to $f(\mathbf{A})\mathbf{b}$ is $\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})=\mathbf{Q}f(\mathbf{T})\mathbf{Q}^{\mathsf{T}}\mathbf{b}$. Define
@@
Proving Theorem 2.1 amounts to proving an upper bound on $\|\mathsf{err}_{k}(r)\|_{2}$. Finally, for any symmetric positive definite matrix $\mathbf{M}$, define
@@
2.3 Simplifying $\mathsf{lan}_{k}(r_{j})$ and $\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})$
We begin with a few standard results that will be useful for the proof of Theorem 2.1.
Lemma 2.2 ([DK89, Saa92]).
@@ 2.3 Simplifying $\mathsf{lan}_{k}(r_{j})$ and $\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})$
Using that $\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{Q}=\mathbf{T}$, we find that $\mathbf{A}^{j}\mathbf{b}=\mathbf{Q}\mathbf{T}^{j}\mathbf{Q}^{\mathsf{T}}\mathbf{b}$, and thus, by linearity, $n(\mathbf{A})\mathbf{b}=\mathbf{Q}n(\mathbf{T})\mathbf{Q}^{\mathsf{T}}\mathbf{b}=\mathsf{lan}_{k}(n)$.
Lemma 2.4.
Let $\mathbf{A}_{j}$ be as in Eq. 9. Then, for any $k\geq 0$, $\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})=\mathbf{Q}(\mathbf{T}-z_{j}\mathbf{I})^{-1}\mathbf{Q}^{\mathsf{T}}r_{j-1}(\mathbf{A})\mathbf{b}$.
Proof 2.5.
@@
Recall that $\mathbf{A}_{j}=\pm(\mathbf{A}-z_{j}\mathbf{I})$. Since $\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{Q}=\mathbf{T}$ and $\mathbf{Q}^{\mathsf{T}}\mathbf{Q}=\mathbf{I}$, we have $\mathbf{Q}^{\mathsf{T}}\mathbf{A}_{j}\mathbf{Q}=\pm(\mathbf{Q}^{\mathsf{T}}(\mathbf{A}-z_{j}\mathbf{I})\mathbf{Q})=\pm(\mathbf{T}-z_{j}\mathbf{I})$. Noting also that $\mathbf{A}_{j}r_{j}(\mathbf{A})=\pm r_{j-1}(\mathbf{A})$ and plugging into Eq. 10 gives the result.
2.4 A telescoping sum for the error in terms of $\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})$
@@
Focusing on the second term and using Lemma 2.4,
@@
Substituting this into Eq. 11, we have
@@
Next, notice that $\mathsf{opt}_{k}(r_{1},\mathbf{A}_{1})$ is exactly the Lanczos approximation to $r_{1}(\mathbf{A})\mathbf{b}$. Indeed, since $\mathbf{Q}^{\mathsf{T}}\mathbf{Q}=\mathbf{I}$, Lemmas 2.2 and 2.4 imply
@@
Using this as a base case, we can repeatedly apply our decomposition of $\mathsf{err}_{k}(r_{j})$ in Eq. 12 to obtain
@@
2.5 Bounding each term in the telescoping sum and combining
@@
Combining these two steps proves the lemma.
With the above results in place, we can prove Theorem 2.1. We will apply the triangle inequality to the telescoping sum for $\mathsf{err}_{k}(r)$ (Lemma 2.6) and bound each term by the error of the optimal approximation to $r(\mathbf{A})\mathbf{b}$ (Lemma 2.8).
Proof 2.10 (Proof of Theorem 2.1).
Focus on a single term in the sum in Lemma 2.6. First we will get rid of $\mathbf{Q}$ and $\mathbf{Q^{\mathsf{T}}}$. For $1\leq j<q$, using that $\mathbf{Q}^{\mathsf{T}}\mathbf{Q}=\mathbf{I}$ we have that
@@
Also note that by the convention adopted in Lemma 2.6, for $j=q$, $\prod_{i=j+1}^{q}\mathbf{Q}(\mathbf{T}-z_{k}\mathbf{I})^{-1}\mathbf{Q}^{\mathsf{T}}=\mathbf{I}=m_{j+1,q}(\mathbf{T})$. Next we claim that
@@
We now apply the triangle inequality to the telescoping sum of Lemma 2.6:
@@
Applying Eq. 13 and Eq. 14 we then have
@@
Finally, using Lemma 2.8 we find
@@
The result follows by plugging into (15).
2.6 Comparison to Lanczos-OR
In [CGMM23] an algorithm called the Lanczos method for optimal rational function approximation (Lanczos-OR) was developed. For rational functions of the form Eq. 8, when run for $k>\deg(n)$ iterations, Lanczos-OR produces iterates
This implies a 2-norm instance optimality guarantee (Definition 1.1):
This guarantee is similar to and often somewhat better than Theorem 2.1. For example, when $r(x)={1}/{m(x)}$, $\sqrt{\kappa(r(\mathbf{A}))}=\sqrt{\kappa(\mathbf{A}_{1})\cdot\kappa(\mathbf{A}_{2})\cdots\kappa(\mathbf{A}_{q})}$. In this case Eq. 16 improves on Theorem 2.1 by a $q$ factor and a square root. However, the bound is for a different algorithm; it does not extend to the ubiquitous Lanczos-FA method.
Moreover, despite this bound, Lanczos-OR usually performs worse than Lanczos-FA in practice if error is measured in the 2-norm. It is not hard to find examples where Eq. 16 is essentially sharp; see for instance Fig. 6. On the other hand, we have been unable to find any numerical examples where Theorem 2.1 is sharp, suggesting that it may be possible to prove a tighter bound for Lanczos-FA.
2.7 Implications for non-rational functions
In this section, we briefly discuss the implications of instance optimality guarantees for rational functions (like Theorem 2.1) to the problem of deriving guarantees for other (non-rational) functions. In particular, consider any function $f(x)$ that is uniformly well approximated by a low-degree rational function $r(x)$ on $\mathcal{I}=[\lambda_{\textup{min}},\lambda_{\textup{max}}]$, the field of values of $\mathbf{A}$. Furthermore, assume we have the following instance optimality guarantee for rational functions as in Definition 1.1:
We write the constants with subscripts to make it clear that they depend on the choice of approximant $r$. Theorem 2.1 is of this form, though as discussed above, its parameters may be loose. A natural way to approximate $f(\mathbf{A})\mathbf{b}$ would be to fix $r$, a rational approximation to $f$, and then apply Lanczos-FA to $r(\mathbf{A})\mathbf{b}$. However, the following analysis shows that Lanczos-FA applied to $f$ itself is nearly as good.
We can bound the error of Lanczos-FA on $f$ using triangle inequality as follows:
@@ 2.7 Implications for non-rational functions
The first and third terms of Subsection 2.7 are controlled by the maximum error of approximating $f$ with $r$ over $\mathcal{I}$. Specifically, if we let $\|r-f\|_{\mathcal{I}}:=\max_{x\in\mathcal{I}}|r(x)-f(x)|$,
@@
The second term of Subsection 2.7 can controlled using Eq. 17 and a triangle inequality:
@@
Compare the form of this bound to that of Definitions 1.1 and 1.3. It is close to an instance near-optimality guarantee, expect for the first term, which requires $f$ to be uniformly well-approximated by a rational function $r$ on $\mathcal{I}=[\lambda_{\textup{min}},\lambda_{\textup{max}}]$. This is still much stronger than FOV optimality (Fact 1), which requires $f$ to be uniformly well-approximated by a polynomial to guarantee that Lanczos-FA provides a good approximation. There are many functions with lower degree rational approximations than polynomial approximations, even when we require the rational function $r$ to have poles only in $\mathbb{R}\setminus\mathcal{I}$ (as in our Theorem 2.1). For example, a uniform polynomial approximation to the square root on a strictly positive interval $[\lambda_{\textup{min}},\lambda_{\textup{max}}]$ requires degree $\Omega(\sqrt{\lambda_{\textup{max}}/\lambda_{\textup{min}}})$ [Tre19, Chapter 8]. On the other hand, a uniform rational approximation can be obtained with degree only $\Omega(\log(\lambda_{\textup{max}}/\lambda_{\textup{min}}))$ [HHT08, Tre19, Chapter 25]. Likewise, a uniform polynomial approximation to $\exp(-x)$ on the interval $[0,B]$ requires degree $\Omega(\sqrt{B})$ [AA22], but uniform rational approximations can be constructed with no dependence on $B$ at all [SSV75]. For such functions, we expect Eq. 19 to be stronger than Fact 1.
Notice also that, while we assume $f$ is well-approximated by a rational function, we are not required to actually construct the approximation. Indeed, since it holds for any $r$, instead of fixing a rational approximation of a certain degree, Eq. 19 automatically balances $\|r-f\|_{\mathcal{I}}$, which decreases as the degree grows, with $C_{r}$, which may increase as the degree grows. Currently, the instance optimality bound of Theorem 2.1 has a large $C_{r}$ that can grow exponentially with the degree of $r$. However, we conjecture that this dependence can be improved.
3 Spectrum optimality for $\mathbf{A}^{\pm 1/2}\mathbf{b}$
In this section, we prove spectrum optimality results for two especially important matrix functions, square root and inverse square root. We start by stating our result for the later.
Theorem 3.1.
That is, Lanczos-FA applied to the inverse square root satisfies Definition 1.2 with
$C=\frac{3}{\sqrt{\pi k}}\kappa(\mathbf{A}),\qquad c=\frac{1}{2}.$
The proof of Theorem 3.1 relies on comparing the error of the $k$th Lanczos iterate for $x^{-1/2}$ to that of the Lanczos iterate for $x^{-1}$. First, applying a bound from [CGMM22], we use the Cauchy integral formula to upper bound the error of Lanczos-FA on $x^{\pm 1/2}$ by its error on $x^{-1}$ (Lemma 3.3). Second, as Equation Eq. 6 shows, Lanczos-FA is nearly instance optimal for the function $x^{-1}$; that is, it outputs $p(\mathbf{A})\mathbf{b}$ where $p$ is (nearly) the degree $k$ polynomial that best approximates $x^{-1}$. Third, the best degree $k$ polynomial approximation to $x^{-1}$ must have lower error than the best degree $k/2$ approximation to $x^{-1/2}$. This is because any degree $k/2$ approximation to $x^{-1/2}$ can be squared to yield a good degree $k$ polynomial approximation to $x^{-1}$. Combining these three steps upper bounds the error of the $k$th Lanczos-FA iterate for $x^{-1/2}$ by the error of the best degree $k/2$ polynomial approximation of $x^{-1/2}$.
Nearly the same argument can be used to prove spectrum optimality of Lanczos-FA for the function $\mathbf{A}^{-1/n}\mathbf{b}$ for any $n\in\mathbb{N}$ with $C=(2^{n}-1)\cdot\kappa(\mathbf{A})/\sqrt{\pi k}$. Furthermore, recall that using Lemma 1.4, we can convert Theorem 3.1 into a near instance optimality guarantee at the price of strong dependence of $C$ on $\mathbf{b}$.
We next state our optimality result for the matrix square root.
Theorem 3.2.
This bound resembles Definition 1.2 with
$C=\frac{3\kappa(\mathbf{A})^{2}}{k^{3/2}},\qquad c=\frac{1}{2}.$
The proof of Theorem 3.2 is nearly the same as that of Theorem 3.1. Ideally, if $p$ is a polynomial approximation to $x^{1/2}$, we would like to claim that $(p(x)/x)^{2}$ yields a good polynomial approximation to $x^{-1}$. However, since this function is not necessarily a polynomial, we must instead use use $\left(\frac{p(x)-p(0)}{x}\right)^{2}$, which introduces the need to include $\{0\}$ in the maximization on the right-hand side.
Figure 3: The bounds of Theorems 3.1 and 3.2 capture the convergence behavior of Lanczos-FA for $\mathbf{A}^{\pm 1/2}\mathbf{b}$. In particular, they can be tighter than the FOV optimality bound of Fact 1. The predicted rate of convergence is about half that observed for Lanczos-FA in this example; this is due to the parameter $c=1/2$ in both bounds.
3.1 Proofs
We begin by quoting bounds from [CGMM22].
Lemma 3.3.
For all $k\geq 1$, the Lanczos-FA iterate satisfies the bounds

### v3

@@ A.1 Notation
Note that $m_{1,q}=m$, so $r_{q}=r(x)$. Recall that for any function $f(x)$, the Lanczos-FA approximation to $f(\mathbf{A})\mathbf{b}$ is $\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})=\mathbf{Q}f(\mathbf{T})\mathbf{Q}^{\mathsf{T}}\mathbf{b}$. Define
@@
Proving Theorem 4 amounts to proving an upper bound on $\|\mathsf{err}_{k}(r)\|_{2}$. Finally, for any symmetric positive definite matrix $\mathbf{M}$, define
@@
A.2 Simplifying $\mathsf{lan}_{k}(r_{j})$ and $\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})$
We begin with a few standard results that will be useful for the proof of Theorem 4.
Lemma 8 ([druskin_knizhnerman_89, saad_92]).
@@ A.2 Simplifying $\mathsf{lan}_{k}(r_{j})$ and $\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})$
Using that $\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{Q}=\mathbf{T}$, we find that $\mathbf{A}^{j}\mathbf{b}=\mathbf{Q}\mathbf{T}^{j}\mathbf{Q}^{\mathsf{T}}\mathbf{b}$, and thus, by linearity, $n(\mathbf{A})\mathbf{b}=\mathbf{Q}n(\mathbf{T})\mathbf{Q}^{\mathsf{T}}\mathbf{b}=\mathsf{lan}_{k}(n)$. ∎
Lemma 9.
Let $\mathbf{A}_{j}$ be as in Equation 5. Then, for any $k\geq 0$, $\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})=\mathbf{Q}(\mathbf{T}-z_{j}\mathbf{I})^{-1}\mathbf{Q}^{\mathsf{T}}r_{j-1}(\mathbf{A})\mathbf{b}$.
Proof.
@@
Recall that $\mathbf{A}_{j}=\pm(\mathbf{A}-z_{j}\mathbf{I})$. Since $\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{Q}=\mathbf{T}$ and $\mathbf{Q}^{\mathsf{T}}\mathbf{Q}=\mathbf{I}$, we have $\mathbf{Q}^{\mathsf{T}}\mathbf{A}_{j}\mathbf{Q}=\pm(\mathbf{Q}^{\mathsf{T}}(\mathbf{A}-z_{j}\mathbf{I})\mathbf{Q})=\pm(\mathbf{T}-z_{j}\mathbf{I})$. Noting also that $\mathbf{A}_{j}r_{j}(\mathbf{A})=\pm r_{j-1}(\mathbf{A})$ and plugging into Equation 9 gives the result. ∎
A.3 A telescoping sum for the error in terms of $\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})$
@@
Focusing on the second term and using Lemma 9,
@@
Substituting this into Equation 10, we have
@@
Next, notice that $\mathsf{opt}_{k}(r_{1},\mathbf{A}_{1})$ is exactly the Lanczos approximation to $r_{1}(\mathbf{A})\mathbf{b}$. Indeed, since $\mathbf{Q}^{\mathsf{T}}\mathbf{Q}=\mathbf{I}$, Lemmas 8 and 9 imply
@@
Using this as a base case, we can repeatedly apply our decomposition of $\mathsf{err}_{k}(r_{j})$ in Equation 11 to obtain
@@
A.4 Bounding each term in the telescoping sum and combining
@@
Combining these two steps proves the lemma. ∎
With the above results in place, we can prove Theorem 4. We will apply the triangle inequality to the telescoping sum for $\mathsf{err}_{k}(r)$ (Lemma 10) and bound each term by the error of the optimal approximation to $r(\mathbf{A})\mathbf{b}$ (Lemma 11).
Proof of Theorem 4.
Focus on a single term in the sum in Lemma 10. First we will get rid of $\mathbf{Q}$ and $\mathbf{Q^{\mathsf{T}}}$. For $1\leq j<q$, using that $\mathbf{Q}^{\mathsf{T}}\mathbf{Q}=\mathbf{I}$ we have that
@@
Also note that by the convention adopted in Lemma 10, for $j=q$, $\prod_{i=j+1}^{q}\mathbf{Q}(\mathbf{T}-z_{k}\mathbf{I})^{-1}\mathbf{Q}^{\mathsf{T}}=\mathbf{I}=m_{j+1,q}(\mathbf{T})$. Next we claim that
@@
We now apply the triangle inequality to the telescoping sum of Lemma 10:
@@
Applying Equation 12 and Equation 13 we then have
@@
Finally, using Lemma 11 we find
@@
The result follows by plugging into (14). ∎
A.5 Proof of Lemma 5
We can bound the error of Lanczos-FA on $f(x)$ using triangle inequality as follows:
@@ A.5 Proof of Lemma 5
The first and third terms of Section A.5 are controlled by the maximum error of approximating $f(x)$ with $r(x)$ over $\mathcal{I}$. Specifically, if we let $\|r-f\|_{\mathcal{I}}:=\max_{x\in\mathcal{I}}|r(x)-f(x)|$,
@@
The second term of Section A.5 can controlled using Equation 7 and a triangle inequality:
@@
Appendix B Comparison to Prior Work
B.1 Details of existing near-optimality guarantees for Lanczos-FA
In this section, we review in detail the prior analyses of Lanczos-FA cited in Section 1.3. These are the only near-optimality guarantees for Lanczos-FA of which we are aware. Instance optimality trivially holds when $f(x)$ is a polynomial of degree $<k$. In this case, it is well known that $\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})$ exactly applies $f(\mathbf{A})\mathbf{b}$; i.e., $\|f(\mathbf{A})\mathbf{b}-\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}=0$ [saad_92]. When $f(x)=1/x$ and $\mathbf{A}$ is positive definite, the Lanczos-FA algorithm is mathematically equivalent to the well-known conjugate gradient algorithm for solving a system $\mathbf{A}\mathbf{x}=\mathbf{b}$. This implies that Lanczos-FA is the optimal approximation in the Krylov subspace with respect to the $\mathbf{A}$-norm; that is,
where $\kappa(\mathbf{A})$ is the condition number $\lambda_{\textup{max}}/\lambda_{\textup{min}}$ of $\mathbf{A}$. That is, Definition 2 is satisfied with $C=\sqrt{\kappa(\mathbf{A})}$ and $c=1$.
Note that the best polynomial approximation to $1/x$ on $\mathcal{I}=[\lambda_{\textup{min}},\lambda_{\textup{max}}]$ already converges at a geometric rate [jokar_mehri_05]:
That is, Fact 3 suffices to prove the exponential convergence of the conjugate gradient method. However, conjugate gradient (and equivalently Lanczos-FA) often converges even faster, which is explained theoretically by the stronger instance- and spectrum-optimality guarantees that the method satisfies [beckermann_kuijlaars_01, liesen_strakos_13, carson_liesen_strakos_24].
If $\mathbf{A}$ is not positive definite, $\mathbf{T}$ may have an eigenvalue at or near to zero and the error of the Lanczos-FA approximation to $\mathbf{A}^{-1}\mathbf{b}$ can be arbitrarily large. In fact, the same is true for any function $f(x)$ which is much larger on $\mathcal{I}=[\lambda_{\textup{min}},\lambda_{\textup{max}}]$ than on $\Lambda$, the set of eigenvalues of $\mathbf{A}$. However, while the Lanczos-FA iterates may be bad at some iterations, the overall convergence of Lanczos-FA with $f(x)=1/x$ is actually good [chen_meurant_24]. In particular, the convergence of Lanczos-FA can be related to the MINRES algorithm [cullum_greenbaum_96, meurant_tebbens_20], which produces an optimal approximation to $\mathbf{A}^{-1}\mathbf{b}$ with respect to the $\mathbf{A}^{2}$-norm. This allows certain optimality guarantees for MINRES to be transferred to Lanczos-FA, even on indefinite systems. In particular, [chen_meurant_24] asserts that for every $k$, there exists $k^{*}\leq k$ such that [^4] For indefinite matrices, $\kappa(\mathbf{A})$ denotes the ratio of the largest eigenvalue magnitude to the smallest eigenvalue magnitude.
$\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{lan}_{k^{*}}(1/x;\mathbf{A},\mathbf{b})\|_{2}\leq\kappa(\mathbf{A})\sqrt{k+1}\min_{\deg(p)<k}\|\mathbf{A}^{-1}\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}.$ (18)
While (18) does not quite fit the form of Definition 2 because of the $k^{*}$ on the left side, it is similar in spirit. Also note that the dependence on $k$ in the prefactor $\sqrt{k+1}$ is not of great significance. Indeed, the convergence of the optimal polynomial approximation to $1/x$ on two intervals bounded away from zero is geometric.
Finally, a guarantee for the matrix exponential is proved in [druskin_greenbaum_knizhnerman_98, (45)]. They show that the Lanczos-FA iterate satisfies the guarantee
Again, this bound does not quite fit Definition 2 due to the maximization over $s$. The authors of [druskin_greenbaum_knizhnerman_98] state, “It is not known whether the maximum over $s\in[0,t]$ in (45) can be omitted and $s$ set equal to $t$ in the right-hand side.”
B.2 Comparison to Lanczos-OR
In [chen_greenbaum_musco_musco_23] an algorithm called the Lanczos method for optimal rational function approximation (Lanczos-OR) was developed. For rational functions of the form Equation 4, when run for $k>\deg(n)$ iterations, Lanczos-OR produces iterates
This implies a 2-norm instance-optimality guarantee (Definition 2):
This guarantee is similar to and often somewhat better than Theorem 4. For example, when $r(x)={1}/{m(x)}$, $\sqrt{\kappa(r(\mathbf{A}))}=\sqrt{\kappa(\mathbf{A}_{1})\cdot\kappa(\mathbf{A}_{2})\cdots\kappa(\mathbf{A}_{q})}$. In this case Equation 19 improves on Theorem 4 by a $q$ factor and a square root. However, the bound is for a different algorithm; it does not extend to the ubiquitous Lanczos-FA method.
Moreover, despite this bound, Lanczos-OR usually performs worse than Lanczos-FA in practice if error is measured in the 2-norm. It is not hard to find examples where Equation 19 is essentially sharp; see for instance Figure 8. On the other hand, we have been unable to find any numerical examples where Theorem 4 is sharp, suggesting that it may be possible to prove a tighter bound for Lanczos-FA.
Appendix C Instance, spectrum, and FOV optimality
The convergence of Lanczos-FA is entirely determined by spectral properties of $\mathbf{A}$. Therefore, following the classical analysis of $f(x)=1/x$, we can relax the notion of instance optimality (Definition 2) by removing its dependence on $\mathbf{b}$:
Definition 12 (Near Spectrum Optimality).
For an input instance $(f,\mathbf{A},\mathbf{b},k)$, we say that a Krylov method is nearly spectrum optimal with parameters $C,c$ if
$\|f(\mathbf{A})\mathbf{b}-\textsf{{alg}}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}\leq C\|\mathbf{b}\|_{2}\min_{\deg(p)<ck}\|f(\mathbf{A})-p(\mathbf{A})\|_{2}.$
Note that Definition 12 depends only on the spectrum of $\mathbf{A}$, and not on the interaction of the eigenvectors of $\mathbf{A}$ with the starting vector $\mathbf{b}$. Hence, we call it “spectrum optimality”. Note also that the guarantee of Theorem 6 fits this form. We can further relax this notion of optimality by weakening its dependence on $\mathbf{A}$:
Definition 13 (Near FOV Optimality).
For an input instance $(f,\mathbf{A},\mathbf{b},k)$, we say that a Krylov method is nearly FOV optimal with parameters $C,c$ if
Note that Fact 3 fits this form. Each type of optimality tightly relaxes the previous one. Definition 2 implies Definition 12 and, as we show in Theorem 14, for any $f,k$ and $\mathbf{A}$, there exists a “worst-case” choice of $\mathbf{b}$ for which their bounds coincide; likewise, Definition 12 implies Definition 13, and for any $f(x)$ and $k$ there exists a “worst-case” choice of $\mathbf{A}$ and $\mathbf{b}$ for which their bounds coincide.
Theorem 14.